# references.bib

### v1

@@ IV-B Transmitter and Receiver Access Point Selection
The proposed method to minimize $P_{D}$ involves a minimization of the MI w.r.t. APs being ISAC-signal receivers or transmitters.
Consider the set of all APs (initially considered as transmitters), $\mathcal{A}=$$\{a_{l}\}_{1}^{N_{\text{AP}}}$ where $a_{l}$ is each AP $l=1,2,...,N_{AP}$, and $N_{AP}$ is the total number of APs. The problem formulation then becomes
@@
where $\mathcal{R}$ is the set of sensing receiver APs.
@@
$I(y_{i};\mathbf{x}^{s}_{l})\leq\log_{2}\left[\det\left(\mathbf{I}+\mathbf{H}_{i,l}\mathbf{x}^{s}_{l}(\mathbf{x}^{s}_{l})^{H}\mathbf{H}_{i,l}^{H}\right)\right],$ (19)
@@
=$\mathcal{R}=\left\{a_{l}\in\mathcal{A}\,\,\forall l\in\mathbf{l}_{\text{Rx}}:\operatorname*{argmax}_{\mathbf{l}_{\text{Rx}}}\sum\limits_{l\in\mathbf{l}_{\text{Rx}}}\sum\limits_{i=1}^{N_{\text{UE}}}I(y_{i};\mathbf{x}^{s}_{l})\right\}$ (20)
@@ V Results
Results presented in this section are produced using the parameters in Table I unless specified. The system is randomized, with certain constraints [^1] https://gitlab.liu.se/henak56/privacy-for-isac, for each of the $500$ trials used for each data point in all plots.
@@
Path Loss Exponent $2.8$ $P_{max}$ $40$ dBm
Max EM Iterations $100$ $\gamma$ $3$ dB
Max CCP Iterations $100$ $\sigma_{H}$ $-40$ dB
EM Convergence Threshold $1\cdot 10^{-10}$ $\alpha_{j,r}$ $10$ dB
CCP Convergence Threshold $0.1$ $\sigma_{n}$ $-94$ dBm
$N_{\text{AP}}$ $5$ $\eta$ $1$
@@
Figure 2: Probability of target detection, $P_{D}$, with varying min power constraint.
Figure 3: Probability of detection, $P_{D}$, for varying number of receiver APs.
Figure 2 shows $P_{D}$ when using no MI minimization, versus using the framework proposed in this paper. As expected, $P_{D}$ is greater when more power is available to the system, due to the sensing beams being more prevalent within the transmitter beampatterns. Moreover, when the framework is limited to optimizing w.r.t. only one receiver AP, a ceiling in reduction of $P_{D}$ seems to be reached, independent of alterations of the maximum power usage constraint.
As seen in 3, $P_{D}$ decreases with more receiver APs, as there are less ISAC transmitters to triangulate an estimated target position. Furthermore, the difference between the two cases increases as more APs are acting as receivers, which demonstrates a privacy-preservation gain from having more degrees of freedom with the developed method. It is worthwhile to mention that, in both cases, the communication SINR and maximum power consumption constraints are met.
@@ VI Conclusions
Attempting to reduce the probability of detection of a target within a ISAC system, from an adversarial user of the system, was the aim of this paper. Using the proposed MI minimization technique, the detection probability stayed the same or decreased in all cases, but further investigations on the trade-offs between computational efforts and allowing for lower sensing SINR are still to be carried out.

### v2

@@ IV-B Transmitter and Receiver Access Point Selection
The proposed method to counteract the adversarial user, involves a minimization of the MI w.r.t. APs being ISAC-signal receivers or transmitters. Consider the set of all APs (considered as transmitters), $\mathcal{A}=$$\{a_{l}\}_{1}^{N_{\text{AP}}}$ where $a_{l}$ is each AP $l=1,2,...,N_{AP}$, and $N_{AP}$ is the total number of APs. The problem formulation then becomes
@@
where $\mathcal{R}$ is the set of receiver APs, and $\mathbf{x}^{s}_{l}\in\mathbb{C}^{M\times N}$ is the precoded sensing symbol from each AP $l$, i.e. $\mathbf{x}^{s}_{l}=\mathbf{w}_{l,t}\,\mathbf{s}_{s}^{H}$, for $\mathbf{s}_{s}$ being an $N\times 1$ vector of sensing symbols.
@@
$\begin{split}I(y_{i};\mathbf{x}^{s}_{l})&\leq\log_{2}\left(1+\mathbf{h}_{l,i}^{H}\mathbf{x}^{s}_{l}(\mathbf{x}^{s}_{l})^{H}\mathbf{h}_{l,i}\right)\\ &=\log_{2}\left(1+\left|\left|(\mathbf{x}^{s}_{l})^{H}\mathbf{h}_{l,i}\right|\right|^{2}\right),\end{split}$ (19)
where $\mathbf{h}_{l,i}\in\mathbb{C}^{M}$ is the channel vector between AP $l$ and user $i$. Exploiting the monotonic properties of the $\log_{2}$ function, in conjunction with (19), enables the minimization problem (18) to be rewritten as
$\min_{\mathcal{R}\subset\mathcal{A}}\quad\sum\limits_{i=1}^{N_{\text{UE}}}\sum\limits_{l=1}^{N_{\text{AP}}}\left|\left|(\mathbf{x}^{s}_{l})^{H}\mathbf{h}_{l,i}\right|\right|^{2}.$ (20)
@@
=$\mathcal{R}=\left\{a_{l}\in\mathcal{A}\,\,\forall l\in\mathbf{l}_{\text{Rx}}:\operatorname*{argmax}_{\mathbf{l}_{\text{Rx}}}\sum\limits_{l\in\mathbf{l}_{\text{Rx}}}\sum\limits_{i=1}^{N_{\text{UE}}}\left|\left|(\mathbf{x}^{s}_{l})^{H}\mathbf{h}_{l,i}\right|\right|^{2}\right\}.$ (21)
Equivalently, defining a matrix whose columns are the simplified MI contributions from one AP to all users (and thus the rows are the MI contributions from all APs to one user) as
$\mathbf{M}=\begin{bmatrix}\left|\left|(\mathbf{x}^{s}_{1})^{H}\mathbf{h}_{1,1}\right|\right|^{2}&\dots&\left|\left|(\mathbf{x}^{s}_{N_{\text{AP}}})^{H}\mathbf{h}_{N_{\text{AP}},1}\right|\right|^{2}\\ \vdots&\ddots&\vdots\\ \left|\left|(\mathbf{x}^{s}_{1})^{H}\mathbf{h}_{1,N_{\text{UE}}}\right|\right|^{2}&\dots&\left|\left|(\mathbf{x}^{s}_{N_{\text{AP}}})^{H}\mathbf{h}_{N_{\text{AP}},N_{\text{UE}}}\right|\right|^{2}\end{bmatrix},$
facilitates a modification of the solution into
=$\mathcal{R}=\left\{a_{l}\in\mathcal{A}\,\,\forall l\in\mathbf{l}_{\text{Rx}}:\operatorname*{argmax}_{\mathbf{l}_{\text{Rx}}}\sum\limits_{l\in\mathbf{l}_{\text{Rx}}}\left|\left|\mathbf{M}_{l}\right|\right|\right\},$ (22)
with $\mathbf{M}_{l}$ being the $l$th column vector of $\mathbf{M}$.
Algorithm 1 summarizes the proposed privacy-preserving framework.
Algorithm 1 Privacy-Preserving Framework
0:  Since the framework is proposed for a centralized control system, general system knowledge is assumed.
1:  $\mathcal{R}_{p=0}\leftarrow\emptyset$
2:  $p\leftarrow 0$
3:  repeat
4:     $p\leftarrow p+1$
5:     Solve (12) for $\mathbf{W}_{p}$ by using (15) and (17), for AP receiver configuration $\mathcal{R}_{p-1}$
6:     Generate sensing signals $\mathbf{x}^{s}_{l}$ for each AP $l$, using precoder matrix $\mathbf{W}_{p}$
7:     Get $\mathcal{R}_{p}$ from (22) by using $\mathbf{x}^{s}_{l}$ and $\mathbf{h}_{l,i}$, assuming perfect channel estimation by the centralized control unit
8:  until $\mathcal{R}_{p}=\mathcal{R}_{p-1}$ or $p\geq 50$
8:  $\mathbf{W}_{p},\mathcal{R}_{p}$
@@ V Results
Results presented in this section are produced [^1] https://github.com/ISAC-group-LiU/Privacy-preservation using the parameters in Table I unless specified. The system is randomized, with certain constraints, for each of the $1000$ trials used for each data point in all plots. The effectiveness of the attack model is measured in probability of detection, $P_{D}$, of the target, by the attacker, and is the amount of successful target position estimations, divided by the total number of trials.
@@
Path Loss Exponent $3$ $N$ $16$
Max EM Iterations $100$ $P_{max}$ $35$ dBm
Max CCP Iterations $100$ $\gamma$ $3$ dB
EM Convergence Threshold $1\cdot 10^{-10}$ $\sigma_{H}$ $-40$ dB
CCP Convergence Threshold $0.1$ $\alpha_{j,r}$ $10$ dB
$N_{\text{AP}}$ $6$ $\sigma_{n}$ $-94$ dBm
$N_{\text{Rx}}$ $1$ $\eta$ $1$
@@
Figure 2: Probability of target detection, with varying min power constraint.
Figure 3: Probability of detection, and sensing SINR, for varying number of receiver APs.
Figure 2 shows $P_{D}$ when using no MI minimization, versus using the framework proposed in this paper, when adjusting the transmit power constraint (12c). As expected, $P_{D}$ is greater when more power is available to the system, due to the sensing beams being more prevalent within the transmitter beampatterns. Moreover, when the framework operates within a system with an abundance of power usage, it performs worse than if the power supply is limited. Again, this is expected considering the sensing beams becoming more difficult to obfuscate when they are stronger.
As seen in Figure 3, $P_{D}$ decreases with more receiver APs, as there are less ISAC transmitters to triangulate an estimated target position. Furthermore, the difference between the two cases increases as more APs are acting as receivers, which demonstrates a privacy-preservation gain from having more degrees of freedom with the developed method. It is worthwhile to mention that, in both cases, the communication SINR and maximum power consumption constraints are met.
Additionally, the effects of the proposed framework on the sensing SINR is shown in Figure 3. The improvement in $P_{D}$ is attained with minimal (to none) sensing SINR loss. Although, the dynamic change on sensing nodes may require increased complexity in the coordination and fusion of sensing information.
@@ VI Conclusions
This paper introduced an attack model in the design of ISAC systems in cell-free MIMO networks. The attacker, a user of the network, leaks sensing information contained in the design of the precoders for ISAC operations. As a countermeasure, we proposed a framework to reduce the probability of detection of the target, within an ISAC system, from the attacker by jointly designing the precoders and selection of sensing units based on mutual information. Our results show that the probability of detection is effectively reduced without significant drawbacks in terms of sensing performance. It is proved that even a simple MI minimization approach can contribute to the prevention of sensing information leakage. While more sophisticated methods can be further analyzed for privacy-preservation on ISAC systems, this paper serves as a benchmark for further development of countermeasures of this type of privacy intrusion.